### v1

@@ 1.2 Proof ideas
We show that if an adaptive query algorithm makes less than $q<\Omega(p^{t/2})$ queries, then the total variation distance $||\mathcal{D}_{Yes}-\mathcal{D}_{No}||_{TV}$ between the two distributions $\mathcal{D}_{Yes}$ and $\mathcal{D}_{No}$ is $\leq\frac{1}{3}$. This proves that any adaptive query algorithm which distinguishes between $\mathcal{D}_{Yes}$ and $\mathcal{D}_{No}$, i.e., where $||\mathcal{D}_{Yes}-\mathcal{D}_{No}||_{TV}>\frac{1}{3}$, must make at least $\Omega(p^{t/2})$ queries. This essentially proves Theorem 1.9.
1.2.1 Comparison with Gopalan et al. [15]
In this section, we discuss the main differences and points between our proof and the one by Gopalan et al. [15]. Our proofs (and the analysis of the testing algorithm) are more elaborate, subtle, case-dependent, and complicated than [15].
Linear dependence is bad. [15] used the similar idea in Theorem F.2. However, their proof took advantage of the fact that any two distinct non-zero vectors in $\mathbb{Z}_{2}^{n}$ are linearly independent. This is not true for $\mathbb{Z}_{p}^{n}$, as distinct vectors can be scalar multiples of each other; e.g., $(1,\cdots,1)$ and $(2,\cdots,2)$ in $\mathbb{Z}_{p}^{n}$, for any $p\geq 3$. Thus, our proof technique is capable of handling and overcoming these difficulties effectively. This is also why our analysis of the structure theorems and algorithms are very much case-dependent.
Different granularity. In case of $\mathbb{Z}_{2}^{n}$,  [15] defined a function $f:\mathbb{Z}_{2}^{n}\to\{-1,+1\}$ to be $k$-granular, where $k$ is a positive integer, if all of its nonzero Fourier coefficients are an integer multiple of $\frac{1}{2^{k}}$. In the case of $\mathbb{Z}_{p}^{n}$, since the Fourier coefficients are truly complex numbers, we define a function $f:\mathbb{Z}_{p}^{n}\to\{-1,+1\}$ to be $k$-granular if all of its nonzero Fourier coefficients are of the form $\frac{g(\omega_{p})}{p^{k}}$ (see Definition 1.1), where $g\in\mathbb{Z}[X]$ and $\omega_{p}$ is a primitive $p^{th}$ root of unity.
Algebraic integer lower bounds are harder. While proving a lower bound on the absolute value of the Fourier coefficients, in the case of $\mathbb{Z}_{2}^{n}$, proving a lower bound on the magnitude of a Fourier coefficient  [15] translates to proving a lower bound on $|g(-1)|/2^{k}$ (since $\omega_{2}=-1$). Note that $|g(-1)|\geq 1$, because $g(-1)\in\mathbb{Z}\setminus\{0\}$; hence the lower bound follows. However, for a general prime $p$, $\omega_{p}$ is truly a complex number, and hence, $g(\omega_{p})$ is no longer an integer (and therefore, there is no way to conclude that $|g(\omega_{p})|\geq 1$). We use a non-trivial result from algebraic number theory (Theorem 2.1) to tackle this problem, in the proof of Theorem 1.5. We also crucially use the fact that the sum of the absolute values of the coefficients of $g$ is at most $p^{k}$, to get an upper bound on the quantities $|g(\omega_{p}^{i})|$, for any $i\in[p-1]$. Clearly, this part requires much more non-triviality than [15].
Do not expect a tight linear lower bound. For $p=2$, it follows from  [15] that the lower bound on the absolute value of the Fourier coefficients is equal to $\frac{1}{s_{f}}$. We show in Theorem 1.6 that there exists a family of Boolean valued functions on $\mathbb{Z}_{p}^{n},\ p>3$ such that $\min_{\chi\in\operatorname{supp}(f)}\left|\widehat{f}(\chi)\right|\leq 1/s_{f}^{1+\alpha_{p}}$, hence proving that the lower bound on the absolute value of the Fourier coefficients is not linear in $\frac{1}{s_{f}}$ in case of $p>3$.
Generalizing the notion of degree. For functions from $\mathbb{Z}_{2}^{n}$ to $\{-1,+1\}$, it is known that $\widehat{f}=\frac{1}{p^{n}}H_{n}f$, where $H_{n}$ denotes the Hadamard matrices (see [7]). Then, the degree $\deg_{2}$ is defined naturally in the case of $\mathbb{Z}_{2}^{n}$. The definition of $\deg_{p}$ becomes slightly challenging, since the relation $\widehat{f}=\frac{1}{p^{n}}H_{n}f$ is simply false over $\mathbb{Z}_{p}^{n}$. To define $\deg_{p}$ for Boolean valued functions on $\mathbb{Z}_{p}^{n}$ (see Definition H.6), where $p$ is an odd prime, we have defined the matrix $V_{1}$ using Vandermonde matrix, and have inductively defined the matrices $V_{n}$, where $n$ is a positive integer, via Kronecker product. Then we have shown that $\widehat{f}=\frac{1}{p^{n}}V_{n}f$, which helps us to define $\deg_{p}(f)$. We claim that this is an appropriate generalization of degree, since $H_{n}=V_{n}$, over $\mathbb{Z}_{2}^{n}$.
Finally, we remark that because of the absence of a linear structure on the solution space of systems of equations over arbitrary Abelian groups (e.g. $\mathbb{Z}_{p^{2}}$), we could not generalize this to arbitrary Abelian groups.
2 Lower bound on the Fourier coefficients
In this section, we will prove Theorem 1.5 assuming Theorem 1.3, which gives us a lower bound on the Fourier coefficients of functions from $\mathbb{Z}_{p}^{n}$ to $\{-1,+1\}$, where $p$ is a prime number. This is a generalization on the granularity of a function $f:\mathbb{Z}_{2}^{n}\to\mathbb{R}$ when the domain of the function is $\mathbb{Z}_{p}^{n}$. We will use the following theorem; for more details see Section A.3.
Theorem 2.1.
[…]
Proof 2.2 (Proof of Theorem 1.5).
If we put $\mu=0$ in Theorem 1.3, we get that there exists a $g\in\mathbb{Z}[X]$, such that $|\widehat{f}(\chi)|\geq|g(\omega_{p})/p^{k}|$, where $k=\lceil\log_{p}s_{f}\rceil+1$, $s_{f}$ being the sparsity of the Boolean valued function $f$. We know by Theorem 2.1 that $|\prod_{i=1}^{p-1}g(\omega_{p}^{i})|\geq 1$. Therefore,
[…]
For $p=2$, $g(\omega_{p}^{i})$ is an integer, so
$\bigg{|}\frac{g(\omega_{p})}{p^{k}}\bigg{|}\;\geq\;\frac{1}{4s_{f}}$ (2.1)
For $p\neq 2$, the conjugate of $g(\omega_{p}^{i})$, namely $\overline{g(\omega_{p}^{i})}$, is nothing but $=g(\omega_{p}^{p-i})$. Since $|g|_{1}\leq p^{k}$, it follows that for any $i\in[p-1]$, $|g(\omega_{p}^{i})/p^{k}|\leq 1$. Therefore,
[…]
$\Rightarrow\prod_{i=1}^{(p-1)/2}\bigg{|}\frac{g(\omega_{p}^{i})}{p^{k}}\bigg{|}^{2}\;\geq\;(\frac{1}{p^{k}})^{p-1}$
$\Rightarrow\bigg{|}\frac{g(\omega_{p})}{p^{k}}\bigg{|}^{2}\;\geq\;\frac{1}{p^{k(p-1)}}$
$\Rightarrow\bigg{|}\frac{g(\omega_{p})}{p^{k}}\bigg{|}\;\geq\;\frac{1}{p^{k(p-1)/2}}\geq\frac{1}{(p^{2}s_{f})^{(p-1)/2}}\;.$ (2.2)
So, from Equation 2.1 and 2.2, we have
$\bigg{|}\frac{g(\omega_{p})}{p^{k}}\bigg{|}\;\geq\;\frac{1}{(p^{2}s_{f})^{\lceil(p-1)/2\rceil}}\;.$
Remark 2.3.
Let $p$ be a prime. The proof-technique of Theorem 1.5 gives the following. If the Fourier coefficients of a Boolean function $f$ are $k$ granular, where $k=\lceil\log_{p}s\rceil+1$, then the Fourier coefficients of $f^{2}$ are $2k$-granular, and their absolute values are $\geq\frac{1}{(p^{2}s)^{(p-1)}}$.
3 Sparse Boolean-valued function with small Fourier coefficients
In this section, for a fixed prime $p\geq 5$, and arbitrarily large $s$, we give an example of a function $f:\mathbb{Z}_{p}^{n}\to\{-1,+1\}$, such that the minimum of the absolute value of its Fourier coefficients is at most $o\left(1/s\right)$. In this section, we give the details of the construction of Theorem 1.6; the details are presented in the Section D.1 and Section D.2.
@@ 3 Sparse Boolean-valued function with small Fourier coefficients
There is a Fourier coefficient of $\operatorname{\mathsf{AT}}$, whose absolute value is $\frac{1}{p^{nc}}$, where $c$ is a constant $>1$.
The proof of Lemma 3.2 is given in Section D.1 due to lack of space.
Now we are ready to prove Theorem 1.6.
Proof 3.3 (Proof of Theorem 1.6).
This directly follows, as we can claim from Lemma 3.2 that there exists a family of functions in $\mathbb{Z}_{p}^{n}$ whose absolute value of the minimum coefficient is not linear in $\frac{1}{\text{sparsity}}$, but actually $=\Omega(\frac{1}{\text{sparsity}^{1+\epsilon_{p}}})$. where $\epsilon_{p}>0$, is a $p$-dependent constant.
Lemma 3.4.
The proof of Lemma 3.4 is given in Section D.2 due to lack of space.
4 Testing Algorithm for Sparsity
Here we present the whole algorithm (Algorithm 1) for sparsity testing over $\mathbb{Z}_{p}^{n}$ (see next page). However, we describe the various steps of the algorithm in Section E.1, including the correctness and the analysis of the algorithm in Section E.1 except for the proofs of two crucial lemmas that are proved in Section E.2 and Section E.3.
Algorithm 1 Test Sparsity
1:  Input: $s,\epsilon$, and query access to $f:\mathbb{Z}_{p}^{n}\to\{-1,+1\}$.
3:  Parameter setting: Set $t:=\lceil 2\log_{p}s+\log_{p}20\rceil+1,\tau:=\min(\frac{\epsilon^{2}}{40p^{t}},\frac{1}{(p^{2}s)^{p-1}})$ $M:=O(\log(p^{t})\cdot\frac{1}{\tau^{2}})$
4:  Choose $v_{1},\dots,v_{t}$ linearly independent vectors uniformly at random from $\mathbb{Z}_{p}^{n}$.
5:  Let $H=\operatorname{Span}\{v_{1},\dots,v_{t}\}$
6:  Pick $(z_{1},x_{1}),\dots,(z_{M},x_{M})$ uniformly and independently from $H\times\mathbb{Z}_{p}^{n}$
7:  Query $f(x_{1}),\dots,f(x_{M})$ and $f(x_{1}-z_{1}),\dots,f(x_{M}-z_{M})$
8:  for For every $r\in\mathbb{Z}_{p}^{n}$ do
10:  end for
11:  if number of $r$ for which the estimate of $wt(r+H^{\perp})$ is $\geq\frac{2\tau}{3}$ is $\leq s$ then
12:     Output YES
13:  else
14:     Output NO
15:  end if
5 Conclusion
Gopalan et al. [15] was the first to study the problem of testing Fourier sparsity of Boolean function over $\mathbb{Z}^{n}_{2}$. Along the way, they were able to drive fundamental properties of Boolean functions over $\mathbb{Z}_{2}^{n}$, like Granularity of the Fourier spectrum, that have found many other applications [4, 8]. In this work, we have extended their results for groups that can be written as $\mathbb{Z}_{p_{1}}^{n_{1}}\times\cdots\times\mathbb{Z}_{p_{t}}^{n_{t}}$.
The most natural question that arises from our work will be to study this problem for (finite) general Abelian groups. Unfortunately, our work does not extend to finite Abelian groups, because of the absence of the component-wise vector space structure. In particular, it would be nice to get a lower bound on the absolute value of non-zero Fourier coefficients of a sparse Boolean-valued function over $\mathbb{Z}_{p^{2}}^{n}$. In any way, our results in this paper should be seen as a first step toward solving the lower bound problem over finite Abelian group.
Finally, we ask whether it is possible to show a better ($p$-dependent) lower bound over $\mathbb{Z}_{p}^{n}$ on the query-complexity of any adaptive sparsity testing algorithm.
Appendix A Preliminaries
A.1 Fourier Analysis over $\mathbb{Z}_{p_{1}}^{n_{1}}\times\dots\times\mathbb{Z}_{p_{t}}^{n_{t}}$
$\mathbb{Z}_{p_{1}}^{n_{1}}\times\dots\times\mathbb{Z}_{p_{t}}^{n_{t}}$ forms a finite Abelian group under addition whose order is $p_{1}^{n_{1}}\cdots p_{t}^{n_{t}}$, where $p_{1},\ldots,p_{t}$ are distinct primes. Its individual components form a vector space, that is, $\mathbb{Z}_{p_{i}}^{n_{i}}$ is a vector space over the field $\mathbb{Z}_{p_{i}}$ for all $i\in\{1,2,\ldots t\}$ whenever $p_{i}$’s are primes. Throughout this paper we will denote $\mathbb{Z}_{p_{1}}^{n_{1}}\times\dots\times\mathbb{Z}_{p_{t}}^{n_{t}}$ by $\mathcal{G}$. So $\mathcal{G}$ is a finite Abelian group with $|\mathcal{G}|=p_{1}^{n_{1}}\cdots p_{t}^{n_{t}}$, where |.||.|| . | denotes the order of $\mathcal{G}$.
We will denote this root of unity by $\omega_{p}$ a $p^{th}$ primitive root of unity, that is $e^{2\pi\iota/p}$.
Let us begine by defining the characters of $\mathcal{G}$.
Definition A.1.
(Character) A character of $\mathcal{G}$ is a homomorphism $\chi:\mathcal{G}\to\mathbb{C}^{\times}$ of $\mathcal{G}$, that is, $\chi$ satisfies the following:
$\chi(x+y)=\chi(x)\chi(y),\ x,y\in\mathcal{G}.$
Equivalently, a character $\chi$ of $\mathcal{G}$ can be defined by
$\chi(x_{1},\ldots,x_{t})=\chi_{r_{1}}(x_{1})\cdots\chi_{r_{t}}(x_{t})=\omega_{p_{1}}^{r_{1}\cdot x_{1}}\cdots\omega_{p_{t}}^{r_{t}\cdot x_{t}},$
where $\chi_{r_{i}}$ is a character of $\mathbb{Z}_{p_{i}}^{n_{i}}$ for each $i$ and is defined by $\chi_{r_{i}}(x_{i})=\omega_{p_{i}}^{r_{i}\cdot x_{i}}$, $x_{i},r_{i}\in\mathbb{Z}_{p_{i}}^{n_{i}}$ for all $i$, $r_{i}\cdot x_{i}=\sum_{j=1}^{n_{i}}r_{i(j)}x_{i(j)}$, $r_{i(j)}$ and $x_{i(j)}$ being the $j^{th}$ component of $r_{i}$ and $x_{i}$ respectively (It follows from the fact that any character $\chi$ of $\mathcal{G}$ can be written as the product of characters of $\mathbb{Z}_{p_{1}}^{n_{1}},\ldots,\mathbb{Z}_{p_{t}}^{n_{t}}$). Let us denote this character by $\chi_{r_{1},\ldots,r_{t}}$.
Now let us look at some properties of characters.
Lemma A.2.
@@ A.1 Fourier Analysis over $\mathbb{Z}_{p_{1}}^{n_{1}}\times\dots\times\mathbb{Z}_{p_{t}}^{n_{t}}$
Definition A.5.
(Fourier transform) For any function $f:\mathcal{G}\to\mathbb{C}$, the Fourier transform $\widehat{f}:\widehat{\mathcal{G}}\to\mathbb{C}$ is defined by
$\widehat{f}(\chi_{r_{1},\ldots,r_{t}})=\frac{1}{|\mathcal{G}|}\sum_{x\in\mathcal{G}}f(x)\omega_{p_{1}}^{-r_{1}\cdot x_{1}}\cdots\omega_{p_{t}}^{-r_{t}\cdot x_{t}},$
where $x_{i},r_{i}\in\mathbb{Z}_{p_{i}}^{n_{i}}$ for all $i$, and $r_{i}\cdot x_{i}=\sum_{j=1}^{n_{i}}r_{i(j)}x_{i(j)}$, $r_{i(j)}$ and $x_{i(j)}$ being the $j^{th}$ component of $r_{i}$ and $x_{i}$ respectively.
Remark A.6.
The Fourier transform of a function $f:\mathcal{G}\to\mathbb{C}$ is defined by
$\widehat{f}(\chi)=\frac{1}{|\mathcal{G}|}\sum_{x\in\mathcal{G}}f(x)\overline{\chi(x)},$
where $\overline{\chi(x)}$ is the conjugate of $\chi(x)$. The Definition A.5 follows from this, as $\chi=\chi_{r_{1},\ldots,r_{t}}$ for some $r_{i}\in\mathbb{Z}_{p_{i}}^{n_{i}},\ i\in\{1,\ldots,t\}$.
@@
$f(x)=\sum_{\chi_{r_{1},\ldots,r_{t}}\in\widehat{\mathcal{G}}}\widehat{f}(\chi_{r_{1},\ldots,r_{t}})\omega_{p_{1}}^{r_{1}\cdot x_{1}}\cdots\omega_{p_{t}}^{r_{t}\cdot x_{t}},$
where $x_{i},r_{i}\in\mathbb{Z}_{p_{i}}^{n_{i}}$ for all $i$, $r_{i}\cdot x_{i}=\sum_{j=1}^{n_{i}}r_{i(j)}x_{i(j)}$, $r_{i(j)}$ and $x_{i(j)}$ being the $j^{th}$ component of $r_{i}$ and $x_{i}$ respectively.
Theorem A.8.
@@
  • The Fourier sparsity $s_{f}$ of a function $f:\mathcal{G}\to\mathbb{C}$ is defined to be the number of non-zero Fourier coefficients in the Fourier expansion of $f$ (Theorem A.7). In this paper, by sparsity of a function, we will mean the Fourier sparsity of the function.

### v2

@@ 1.2 Proof ideas
We show that if an adaptive query algorithm makes less than $q<\Omega(\Delta^{t/2})$ queries, then the total variation distance $\mathcal{D}_{Yes}-\mathcal{D}_{No}||_{TV}$ between the two distributions $\mathcal{D}_{Yes}$ and $\mathcal{D}_{No}$ is $\leq\frac{1}{3}$. This proves that any adaptive query algorithm which distinguishes between $\mathcal{D}_{Yes}$ and $\mathcal{D}_{No}$, i.e., where $\mathcal{D}_{Yes}-\mathcal{D}_{No}||_{TV}>\frac{1}{3}$, must make at least $\Omega(p^{t/2})$ queries. This essentially proves Theorem 1.9.
2 Preliminaries
2.1 Fourier Analysis over $\mathcal{G}$
Any finite Abelian group $\mathcal{G}$ can be written as $\mathbb{Z}_{{p_{1}}^{m_{1}}}\times\dots\times\mathbb{Z}_{{p_{T}}^{m_{T}}}$, where $p_{i},\ i\in[T]$, are primes and not necessarily distinct. Throughout this paper, we will denote $\mathbb{Z}_{{p_{1}}^{m_{1}}}\times\dots\times\mathbb{Z}_{{p_{T}}^{m_{T}}}$ by $\mathcal{G}$. So $\mathcal{G}$ is a finite Abelian group with $|\mathcal{G}|=p_{1}^{m_{1}}\cdots p_{T}^{m_{T}}$, where $|.|$ denotes the order of $\mathcal{G}$. We will denote this root of unity by $\omega_{N}$ a $N^{th}$ primitive root of unity, that is $e^{2\pi\iota/N}$. Let us begin by defining the characters of $\mathcal{G}$.
@@ 2.1 Fourier Analysis over $\mathcal{G}$
(Character) For $\mathcal{G}=\mathbb{Z}_{p_{1}^{m_{1}}}\cdots\mathbb{Z}_{p_{T}^{m_{T}}}$ a character of the group $\mathcal{G}$ is a homomorphism $\chi:\mathcal{G}\to\mathbb{C}^{\times}$ of $\mathcal{G}$, that is, $\chi$ satisfies the following: for all $x,y\in\mathcal{G}$ we have $\chi(x+y)=\chi(x)\chi(y)$.
Equivalently, a character of $\mathcal{G}$ is of the form
$\chi(x^{(1)},\dots,x^{(T)})=\chi_{r^{(1)}}(x^{(1)})\cdots\chi_{r^{(T)}}(x^{(T)}),$
where $r^{(i)}\in\mathbb{Z}_{p_{i}^{m_{i}}}$ and $\chi_{r^{(i)}}$ is a character of $\mathbb{Z}_{p_{i}^{m_{i}}}$ and is defined by $\chi_{r^{(i)}}(x^{(i)})=\omega_{p_{i}^{m_{i}}}^{r^{(i)}x^{(i)}}$.
Thus fo r any $(r^{(1)},\dots,r^{(T)})\in\mathcal{G}$ we can define a corresponding character of $\mathcal{G}$ as $\chi_{r^{(1)},\dots,r^{(T)}}$ that on input $x=(x^{(1)},\dots,x^{(T)})$ is defined as
$\chi_{(r^{(1)},\ldots,r^{(T)})}(x^{(1)},\ldots,x^{(T)})$ $=\omega_{p_{1}^{m_{1}}}^{r^{(1)}\cdot x^{(1)}}\cdots\omega_{p_{T}^{m_{T}}}^{r^{(T)}\cdot x^{(T)}}$ (2.2)
$=\omega_{N}^{\sum_{i=1}^{T}r^{(i)}\cdot x^{(i)}\frac{\Delta}{p_{i}^{m_{i}}}\pmod{\Delta}}$, (2.3)
where $\Delta=\operatorname{LCM}\{p_{1}^{m_{1}},\ldots,p_{T}^{m_{T}}\}$.
Now let us look at some properties of characters.
@@
Definition 2.7.
(Fourier transform) For any $\mathcal{G}=\mathbb{Z}_{p_{1}^{m_{1}}}\times\cdots\times\mathbb{Z}_{p_{T}^{m_{T}}}$ and any function $f:\mathcal{G}\to\mathbb{C}$, the Fourier transform $\widehat{f}:\widehat{\mathcal{G}}\to\mathbb{C}$ is
$\widehat{f}(\chi_{r^{(1)},\ldots,r^{(t)}})=\frac{1}{|\mathcal{G}|}\sum_{x\in\mathcal{G}}f(x)\omega_{p_{1}^{m_{1}}}^{-r^{(1)}\cdot x^{(1)}}\cdots\omega_{p_{T}^{m_{T}}}^{-r^{(T)}\cdot x^{(T)}},$
where $x=(x^{(1)},\dots,x^{(T)})$.
The Fourier transform of a function $f:\mathcal{G}\to\mathbb{C}$ is defined as $\widehat{f}(\chi)=\frac{1}{|\mathcal{G}|}\sum_{x\in\mathcal{G}}f(x)\overline{\chi(x)}$, where $\overline{\chi(x)}$ is the conjugate of $\chi(x)$.
@@
$f(x)=\sum_{\chi_{r^{(1)},\ldots,r^{(T)}}\in\widehat{\mathcal{G}}}\widehat{f}(\chi_{r^{(1)},\ldots,r^{(T)}})\chi_{r^{(1)},\dots,r^{(T)}}(x),$ (2.9)
where $x=(x^{(1)},\dots,x^{(T)})$.
Theorem 2.10.
@@
  • The Fourier sparsity $s_{f}$ of a function $f:\mathcal{G}\to\mathbb{C}$ is defined to be the number of non-zero Fourier coefficients in the Fourier expansion of $f$ (Theorem 2.8). In this paper, by sparsity of a function, we will mean the Fourier sparsity of the function.